OpenStudy (anonymous):

Simplify this complex fraction: (4/x+5)/2

OpenStudy (jhannybean):

Remember \[\frac{(a/b)}{c} = \frac{a}{bc}\]

OpenStudy (jhannybean):

\[\dfrac{\dfrac{4}{x+5}}{2} = \frac{4}{2(x+5)} = \frac{4}{2} \cdot \frac{1}{x+5}\]NOw simplify